OpenStudy (anonymous):

find all value of m such that sqrt(x) = m(x-2) has a solution greater than 4

OpenStudy (dan815):

m>4 right

OpenStudy (empty):

Square both sides of the equation I guess

OpenStudy (dan815):

are u looking for int solutions?

OpenStudy (anonymous):

@dan815 no, real numbers solution

OpenStudy (anonymous):

@Empty I guess you could do so. I've attempted and with a help with wolfram alpha, it got 0 < m < 1. But the arithmetic is just way tedious.

OpenStudy (dan815):

ok well look at the behavior of x/(x-2)^2, find its critical, and inflection points also intersect it with y=4^2 LINE

OpenStudy (anonymous):

just wondering if there is any other easier method

OpenStudy (empty):

Solve for x, then set that value to be greater than 4. Hmmm Ohhh so you're looking for a trick, like a nice way to do it ok that makes this much more fun hahah ok hmmm give me a minute.

OpenStudy (dan815):

oh so x>4 not m

OpenStudy (empty):

Just plug in x=4 and solve for m=1. Then you know m has to be greater than that, cause the square root is always positive, done.

OpenStudy (anonymous):

@dan815 yeah, it's about m that makes x > 4 as solution

OpenStudy (dan815):

you can also think about it like a sqrt function intersecting with a line displaced 2 units right, and m is the slope

OpenStudy (anonymous):

@Empty huhm... so sqrt(4) = m(4-2), which gives m = 1. And then?

OpenStudy (dan815):